OpenStudy (anonymous):

solve using cramers rule please show work x-2y=14 y+2z=11 2x+z=16

OpenStudy (anonymous):

why cramer's lol that rule is so complicated when it is so simple

OpenStudy (anonymous):

I know but I have to learn it

OpenStudy (anonymous):

Okay, can you make the matrix and find the determinant?

OpenStudy (anonymous):

no only cramers

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & -2 & 0 \\ 0 & 1 & 2 \\2 & 0 & 1\end{matrix}\right]\left[\begin{matrix}x \\ y \\z\end{matrix}\right]= \left[\begin{matrix}14 \\ 11 \\16\end{matrix}\right]\] do the process and you will ger x = 4 y = -5 z = 8

OpenStudy (anonymous):

thank you so much may you help me with one more please?